I need the solution to: 3(x-4)=18x
Found 2 solutions by checkley77, stanbon:
Answer by checkley77(9435)

3(x-4)=18x
3x-12=18x
3x-18x=12
-15x=12
x=12/-15
x=-4/5 answer.
proof:
3(-4/5-4)=18*-4/5
3(-4-20)/5=-72/5
3(-24/5=-72/5
-72/5=-72/5

Answer by stanbon(29620)

3(x-4)=18x
3x-12 = 18x
15x = -12
x = -12/15
x = -4/5
============
Cheers,
Stan H.